Factorise. p^4 –81 ????????????????????????????????

Factorise.
p^4 –81 ????????????????????????????????

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

p^4–81
= (p^2)^2-(9)^2
= (p^2-9)(p^2+9)
= (p^2-3^2)(p^2+9)
=(p-3)(p+3)(p^2+9)

Thanks